# Magnetic Field Data Calibration with Transformer Model Using Physical Constraints: A Scalable Method for Satellite Missions, Illustrated by Tianwen-1

### v1

@@ Abstract
This study introduces a novel approach that integrates the magnetic field data correction from the Tianwen-1 Mars mission with a neural network architecture constrained by physical principles derived from Maxwell’s equations. By employing a Transformer-based model capable of efficiently handling sequential data, the method corrects measurement anomalies caused by satellite dynamics, instrument interference, and environmental noise. As a result, it significantly improves both the accuracy and the physical consistency of the calibrated data. Compared to traditional methods that require long data segments and manual intervention often taking weeks or even months to complete this new approach can finish calibration in just minutes to hours, and predictions are made within seconds. This innovation not only accelerates the process of space weather modeling and planetary magnetospheric studies but also provides a robust framework for future planetary exploration and solar wind interaction research.
@@ 1 Introduction
Traditional correction methods rely heavily on human expertise and are rooted in well-established physical and mathematical principles. While these methods have proven effective, they are inherently limited by their long processing times and delays in real-time prediction[7] [6] [4] [2] [1]. In contrast, machine learning models, though rarely applied in this field, offer strong predictive capabilities and the potential for faster computations. This study seeks to address these limitations by combining the strengths of traditional correction methods with the adaptability and efficiency of machine learning models, thereby improving timeliness while ensuring both physical consistency and enhanced real-time performance. This study bridges the gap between data-driven models and physics-based understanding by integrating Maxwell’s equations into the neural network architecture as physical information. The key innovations are:
@@
Neural Network Integration with Physical Information: The integration of physical fields, such as electric and magnetic fields, along with real-time equation calculations, directly into the neural network as input features.
@@ 2.1 Traditional Magnetometer Calibration Methods
Instrument Drift Correction: Instrumental drift occurs due to aging, thermal variations, and operational wear, gradually altering the offset values of magnetometer readings. Correction is achieved by comparing in-flight measurements with known reference conditions, often requiring extensive post-mission data analysis.
Dynamic and Static Field Correction: Dynamic fields generated by onboard spacecraft systems are identified and removed using dual-sensor methods [3], separating spacecraft-induced fields from natural magnetic measurements. Static fields and instrumental offsets are treated together as ”zero offset” values, which require periodic calibration during flight.
@@
Physical Phenomenon-Based Methods:
@@
Segment-Specific Analysis: Data gaps and variability require separate calibration for each continuous data segment, increasing computational and operational overhead.
@@ 2.2.1 Data
Magnetometer data are affected by external factors such as satellite orientation changes and onboard instrument interference. Data from both inner and outer magnetometer probes, as well as their differences, are utilized for correction. Synchronization of data timestamps from multiple sensors, identification and labeling of anomalous data, and resampling at one-minute intervals to reduce noise and computational complexity. We have all the original data from Tianwen-1, including the corrected and published datasets at https://space.ustc.edu.cn/dreams/tw1_momag/?magdata=cal&sr=1.
@@ 2.2.2 Input
The Ampère-Maxwell Law is a fundamental equation in Maxwell’s equations that describes the relationship between the magnetic field , the current density , and the time rate of change of the electric field . Understanding this relationship is crucial for analyzing time-varying electromagnetic fields and the propagation of electromagnetic waves.
@@ 2.2.3 Output
The output data consists of the corrected magnetic field measurements. This data is expected to refine the original measurements by mitigating noise and compensating for any systematic errors, thereby providing more accurate representations of the magnetic fields.
@@ 2.2.4 Neural Network Architecture
The neural network model is a Transformer[5] for time series prediction. It features a custom TransformerBlock that utilizes multi-head attention to capture dependencies across time steps. This block includes a feedforward network with ReLU activation, layer normalization, dropout regularization, and residual connections to enhance training stability and model capacity. The complete architecture comprises an input layer, the TransformerBlock, a flattening layer, a dense layer with ReLU activation, dropout, and an output layer that predicts three target variables. Data preprocessing involves resampling the time series, engineering features by computing derivatives and electric fields, standardizing features and targets, and constructing input sequences. The model is trained using the Adam optimizer with early stopping based on validation loss.
@@ 3 Results
We use two months data for training in 2021. We split the dataset as training set and as test set. Daily comparison plots of actual versus predicted values for demonstrate how our methods effectively predict these components by directly comparing our predicted results with the actual data. Error distribution plots show reduced variance, indicating consistent model performance. We have generated several plots to analyze the performance of our model by comparing our predicted results with the actual data using these methods. Figures 1, 2, 3, 4 and 5 show the comparison of predicted results and actual data, along with the error analysis from our test set.
@@ 4 Discussion
Integration of Physics and Machine Learning: Incorporating physical fields (electric and magnetic) as input features and employing a Transformer-based model enhances prediction accuracy and integrates fundamental physics directly into the machine learning framework.
@@
Scalability and Real-Time Prediction: Our method is adaptable to other satellite missions and can be refined by incorporating additional physical parameters. Moreover, once trained, the model provides near-instantaneous predictions, delivering accurate, high-quality calibrated data that can be readily integrated into mission support and decision-making processes.
@@ 5 Conclusion
This study introduces a scalable, physics-informed machine learning framework that leverages Maxwell’s equations and a Transformer-based neural network to rapidly and accurately correct magnetometer data from the Tianwen-1 Mars mission. By integrating fundamental physical constraints directly into the model, we improve the fidelity and consistency of the corrected magnetic field measurements. Compared to traditional calibration methods that rely on lengthy manual procedures and extended data segments, our approach requires only minutes to hours for training and provides near-instantaneous predictions. The resulting data not only exhibit heightened accuracy and reduced noise but also facilitate more reliable analyses of planetary magnetospheric phenomena and solar wind interactions. Beyond Tianwen-1, this method serves as a flexible and robust blueprint for real-time data calibration in future exploration missions, supporting advanced space weather modeling and significantly enhancing our ability to interpret complex electromagnetic environments in planetary systems.

### v3

@@ Abstract
This study introduces a novel approach that integrates the magnetic field data correction from the Tianwen-1 Mars mission with a neural network architecture constrained by physical principles derived from Maxwell’s equations. By employing a Transformer based model capable of efficiently handling sequential data, the method corrects measurement anomalies caused by satellite dynamics, instrument interference, and environmental noise. In addition to the original approach, we implement two distinct Transformer architectures: a standard Transformer and a physics informed Transformer. The latter incorporates a Fourier Transform branch to extract frequency domain features and enforces a divergence free constraint on the predicted magnetic field, ensuring physical consistency. As a result, while both models significantly improve accuracy and reduce calibration time from weeks or months to minutes or hours. This innovation not only accelerates the process of space weather modeling and planetary magnetospheric studies but also provides a robust framework for future planetary exploration and solar wind interaction research.
@@ 1 Introduction
Traditional correction methods rely heavily on human expertise and are rooted in well established physical and mathematical principles. While these methods have proven effective, they are inherently limited by their long processing times and delays in real time prediction[7] [6] [4] [2] [1]. In contrast, machine learning models, though rarely applied in this field, offer strong predictive capabilities and the potential for faster computations. This study seeks to address these limitations by combining the strengths of traditional correction methods with the adaptability and efficiency of machine learning models, thereby improving timeliness while ensuring both physical consistency and improved real time performance. This study bridges the gap between data driven models and physics based understanding by integrating Maxwell’s equations into the neural network architecture as physical information. The key innovations are:
@@
Neural Network Integration with Physical Information: The integration of physical fields, such as electric and magnetic fields, along with real time equation calculations, directly into the neural network as input features. We introduce two Transformer based architectures. The first is a standard Transformer model for time series prediction, and the second is a physics informed Transformer that integrates a Fourier Transform branch and enforces a divergence free constraint via a dedicated physics layer. These additional components enable the model to capture both time domain and frequency domain features, while simultaneously integrating Maxwell’s equations into the learning process. This dual model strategy bridges the gap between data driven methods and physics based understanding.
@@ 2.1 Traditional Magnetometer Calibration Methods
Instrument Drift Correction: Instrumental drift occurs due to aging, thermal variations, and operational wear, gradually altering the offset values of magnetometer readings. Correction is achieved by comparing in flight measurements with known reference conditions, often requiring extensive post mission data analysis.
Dynamic and Static Field Correction: Dynamic fields generated by onboard spacecraft systems are identified and removed using dual sensor methods [3], separating spacecraft induced fields from natural magnetic measurements. Static fields and instrumental offsets are treated together as ”zero offset” values, which require periodic calibration during flight.
@@
Physical Phenomenon Based Methods:
@@
Segment Specific Analysis: Data gaps and variability require separate calibration for each continuous data segment, increasing computational and operational overhead.
@@ 2.2.1 Data
Magnetometer data are affected by external factors such as satellite orientation changes and onboard instrument interference. Data from both inner and outer magnetometer probes, as well as their differences, are utilized for correction. Synchronization of data timestamps from multiple sensors, identification and labeling of anomalous data, and resampling at one minute intervals to reduce noise and computational complexity. We have all the original data from Tianwen-1, including the corrected and published datasets at https://space.ustc.edu.cn/dreams/tw1_momag/?magdata=cal&sr=1.
@@ 2.2.2 Input
The Ampère-Maxwell Law is a fundamental equation in Maxwell’s equations that describes the relationship between the magnetic field , the current density , and the time rate of change of the electric field . Understanding this relationship is crucial for analyzing time varying electromagnetic fields and the propagation of electromagnetic waves.
@@
We try different combination of inputs, the current inputs give us the best results.
@@ 2.2.3 Output
The output data consists of the corrected magnetic field measurements. This data is expected to refine the original measurements by mitigating noise and compensating for any systematic errors, thereby providing more accurate representations of the magnetic fields. With the introduction of two transformer models, the standard Transformer outputs magnetic field values, while the physics informed Transformer further refines these predictions by integrating electric field computations and enforcing the divergence free condition.
@@ 2.2.4 Neural Network Architecture
Transformer
The neural network model is a Transformer[5] for time series prediction. It features a custom TransformerBlock that utilizes multi-head attention to capture dependencies across time steps. This block includes a feedforward network with ReLU activation, layer normalization, dropout regularization, and residual connections to improve training stability and model capacity. The complete architecture comprises an input layer, the TransformerBlock, a flattening layer, a dense layer with ReLU activation, dropout, and an output layer that predicts three target variables. Data preprocessing involves resampling the time series, engineering features by computing derivatives and electric fields, standardizing features and targets, and constructing input sequences. The model is trained using the Adam optimizer with early stopping based on validation loss.
Physics Informed Transformer with Fourier Transform and Physics Constraint
The second architecture in Figure1 extends the basic Transformer design by incorporating both a Fourier transformation branch and a physics informed constraint. Specifically, a Fourier transformation layer is applied to the entire input sequence to extract frequency domain features, which are then concatenated with the original input data. This augmented input is processed through a custom TransformerBlock that leverages multi-head attention and a feedforward network with dropout and layer normalization. After flattening and passing through additional dense layers, the model outputs predictions for the magnetic field components. Moreover, position features are extracted from the input and used in a PhysicsConstraintLayer to compute the divergence of the predicted magnetic field, enforcing the physical principle of zero divergence in magnetohydrodynamics. The combined loss function during training optimizes both the prediction accuracy and the adherence to the physical constraint.
Fourier Transform Layer Structure
The Fourier Transform layer is designed to extract frequency domain features from the input time series data. The process involves the following steps: The input data is first converted to a complex format to enable complex valued operations. A Fast Fourier Transform is applied to transform the data into its frequency domain representation. The real and the imaginary part of the FFT result is extracted and used as the output of the layer.
Physics Informed Structure
The physics informed structure integrates prior physical knowledge by enforcing a divergence free constraint on the predicted magnetic field. This is achieved through a dedicated layer that computes the divergence residual to ensure that the predicted magnetic field, , satisfies . The procedure is as follows:
1.
Extract the positional features from the input data and combine them with the predicted magnetic field values.
2.
Compute the Jacobian matrix of the predicted magnetic field with respect to the spatial coordinates.
3.
Calculate the divergence by taking the trace of the Jacobian matrix.
4.
Include this divergence as a residual in the loss function to penalize deviations from the physical law.
This physics informed constraint not only improves the model’s accuracy in predicting the magnetic field but also ensures that the predictions adhere to fundamental physical principles.
@@ 3 Results
We use two months data for training in 2021. We split the dataset as training set and as test set. Daily comparison plots of actual versus predicted values for demonstrate how our methods effectively predict these components by directly comparing our predicted results with the actual data. Error distribution plots show reduced variance, indicating consistent model performance. We have generated several plots to analyze the performance of our model by comparing our predicted results with the actual data using these methods. Figures 2, 3, 4, 5 and 6 show the comparison of predicted results and actual data, along with the error analysis from our test set.
The Physics Informed Transformer with Fourier Transform and Physics Constraint achieve a mean absolute error 0.44nT, and has similar results as the Transformer model. Here are the error distributions of the Transformer 7 and The Physics Informed Transformer 8 on the test set.
@@ 4 Discussion
This calibration pipeline effectively reduces noise and anomalies in Tianwen-1 magnetometer data while drastically decreasing the data correction time from days or months to minutes or hours.
@@
Integration of Physics and Machine Learning: Incorporating physical fields (electric and magnetic) as input features and employing a Transformer based model improves prediction accuracy and integrates fundamental physics directly into the machine learning framework.
@@
Scalability and Real Time Prediction: Our method is adaptable to other satellite missions and can be refined by incorporating additional physical parameters. Moreover, once trained, the model provides near-instantaneous predictions, delivering accurate, high quality calibrated data that can be readily integrated into mission support and decision making processes.
@@ 5 Conclusion
This study introduces a scalable, physics informed machine learning framework that leverages Maxwell’s equations and a Transformer based neural network to rapidly and accurately correct magnetometer data from the Tianwen-1 Mars mission. While the original approach based on a standard Transformer already demonstrates significant improvements in both calibration speed and accuracy, our implementation of a second, physics informed Transformer further refines the predictions by integrating Fourier domain insights and enforcing a divergence free condition. By integrating fundamental physical constraints directly into the model, we improve the fidelity and consistency of the corrected magnetic field measurements. Compared to traditional calibration methods that rely on lengthy manual procedures and extended data segments, our approach requires only minutes to hours for training and provides near instantaneous predictions. The resulting data not only exhibit heightened accuracy and reduced noise but also facilitate more reliable analyses of planetary magnetospheric phenomena and solar wind interactions. Beyond Tianwen-1, this method serves as a flexible and robust blueprint for real time data calibration in future exploration missions, supporting advanced space weather modeling and significantly improving our ability to interpret complex electromagnetic environments in planetary systems.